Find the distance between the points (3,17) and (-2,5)
Answer by longjonsilver(2297)

consider the line between the 2 points quoted to be the hypotenuse of a right-angled triangle. Use Pythagoras' theorem to find this distance.
The right angle of your triangle is at (3,5)...sketch it! YOU have to sketch things in maths!!!
horizontal line is length 3-(-2)=5
vertical line is length 17-5 = 12
so, x^2 = 5^2 + 12^2
--> x=13